-12 points DevoreStat9 5.E.072 I have three e ands to take care of in te Administration Building-Let Xi-the time that takes for the rh errand 1-1 2 3 and let X the total time in minutes that I spend walking to and from the building and between each errand. Suppose the x s are independent, and normally distributed it the following means and standard deviations: μ 1-16, σι 4 μ2-7, σ2-1 μ3 10, σ3-2, μ4-13, o4-3. 【 plan to leave my office at precisely 10:00 A M. and wish to post a note on my door that reads, I will retum by t A.M. How long should I estimate my trip will take if I want the probability of the trip taking longer than my estimate to be 0.01? (Round your answer to two decimal places.) My Notes Ask Your Teacher min Need Help?
Answer #1

here total expected time =16+7+10+13=46

std deviation=sqrt(42+12+22+32)=5.477

for top 0.01 level crtiical value =2.33

hence corresponding time =mean+z*std deviaiton =46+2.33*5.477=58.76

( please try 58.74 if this comes wrong)